# Triaxial Magnetars as Sources of Fast Radio Bursts

### v1

@@ Abstract
I suggest that some of the mysterious temporal properties of Fast Radio Bursts (FRB) may be explained if they are produced by dynamically triaxial magnetars. If the bursts are narrowly collimated along open field lines, then observed repeating FRB may be those in which the moment of inertia tensor is only slightly triaxial and the rotation axis, open field lines and radiation point nearly to the observer. Apparently non-repeating FRB may be triaxial with the direction of open field lines and radiation wandering across the sky, reducing their duty factors by several orders of magnitude. A slightly triaxial moment tensor in repeaters moves the line of sight into or out of the radiation pattern or within it, explaining periods of greater or lesser (or absent) activity, and making the probability of detecting a burst and hence the burst rate vary aperiodically. The dynamics of triaxial bodies may also thwart the coherent integration of gravitational signals from fast-rotating accreting neutron stars.
@@ 1 Introduction
In this paper I attribute these phenomena to a triaxial moment of inertia tensor of a magnetically and elastically (by stress frozen-in to the solid component) distorted neutron star. These differences may be enhanced if magnetar-FRB have larger magnetic fields than the known Galactic magnetar-SGR/AXP, implying larger triaxiality of magnetar-FRB moments of inertia.
@@ 2 How Triaxiality May Explain the Three Mysteries?
If a neutron star’s moment of inertia is triaxial by this definition then its rotation is not periodic, but more complex. A triaxial moment of inertia tensor results from frozen-in magnetic stress $B^{2}/8\pi=4\times 10^{28}B_{15}^{2}$ dyne/cm2 and frozen-in elastic stress ${\cal O}(10^{29}$ dyne/cm2) (Horowitz & Kadau, 2009). These may be of similar magnitude and each makes the moment of inertia tensor triaxial. Rotational distortion relaxes to axisymmetry on a hydrodynamic or viscous time scale, is likely smaller than the other distortions for rotational periods $\gtrsim 1\,$s, and does not contribute to rigid body triaxiality.
@@
    Luo et al. (2025) suggested that in repeating FRB the magnetic and spin axes are aligned, while Kumar, Lu & Bhattacharya (2017); Beniamini & Kumar (2025) suggested that FRB radiation is emitted along open field lines. This leads to a possible explanation (Katz, 2025a) of the difference between repeating and apparently non-repeating FRB: Apparently non-repeating FRB have triaxial moment tensors, so that their magnetic poles and directions of radiation wander pseudo-randomly across much of the sky. FRB that are are only minimally tri-axial have their magnetic and rotational axes aligned (dissipation reduces them to their lowest energy state with oblate spheriodal intertia tensors with magnetic, emission and rotation axes aligned); if fortuitously these point to the observer we see them as repeating FRB, while otherwise they are not observable at all.
@@
The first two of these phenomena might be attributed to “plasma weather” (intrinsic stochasticity of emission) in a magnetar’s magnetosphere, but with no explanatory or predictive theory other hypotheses should be considered. “Plasma weather” cannot explain the absence of periodicity (Katz, 2025b) in a rotating neutron star with a spheroidal (axisymmetric) moment of inertia tensor because the observable “weather” would be modulated at the rotation period (as is Solar activity).
@@ 3 Euler Equations
To make order of magnitude estimates, replace the time derivatives by the algebraic factor $1/\tau$, where $\tau$ is a characteristic time scale, the fractions by $\Delta I/I$, a measure of the asymmetry of the moment of inertia tensor, and assume $I_{1}\approx I_{2}\approx I_{3}$ to high accuracy, as the case for a nearly spherical body, but that all the $|I_{i}-I_{j}|$ ($i\neq j$) are of the same order of magnitude.
@@
If the misalignment is comparable to the beam width then the probability of detecting a burst depends on how close to the line of sight is the beam axis because the radiated intensity decreases with increasing deviation. This varies aperiodically if the deviations from an isotropic moment tensor are triaxial (comparable in magnitude, however small). If the observer remains close to the beam axis the probability of detecting a burst depends on how far he is from the axis. This may explain the failure to find periodicity in the rate of detection of bursts in active repeating FRB Katz (2025b).
Eq. 3 may also be applied to spinning accreting neutron stars that have been suggested as sources of detectable gravitational waves. In order to emit gravitational waves they must not be axisymmetric. That need not, in principle, require them to be triaxial; for example, a dynamically prolate spheroid spinning about one of its major axes has an oscillating and radiating gravitational quadrupole moment. However, that geometry is contrived and implausible; such a neutron star would relax to a nearly oblate configuration. If spheroidal, it would not radiate gravitational waves; the deviations (resulting from magnetic or elastic stress) from an oblate spheroid would likely make it triaxial as well as having a rotating gravitational quadrupole. For a fast-spinning neutron star with the previously estimated elastic stress $\tau$ might be of order minutes. This would thwart the coherent integration required to detect a weak periodic gravitational strain.

### v2

@@ Abstract
I suggest that some of the mysterious temporal properties of Fast Radio Bursts (FRB) may be explained if they are produced by dynamically triaxial magnetars. If the bursts are narrowly collimated along open field lines, then observed repeating FRB may be those in which the moment of inertia tensor is only slightly triaxial and the rotation axis, open field lines and radiation point nearly to the observer. Apparently non-repeating FRB may be triaxial with the direction of open field lines and radiation wandering across the sky, reducing their duty factors by several orders of magnitude. A slightly triaxial moment tensor in repeaters moves the line of sight into or out of the radiation pattern or within it, explaining periods of greater or lesser (or absent) activity, and making the probability of detecting a burst and hence the burst rate vary aperiodically. The dynamics of triaxial bodies might thwart the coherent integration of gravitational signals from rotating neutron stars.
@@ 1 Introduction
In this paper I attribute these phenomena to a triaxial moment of inertia tensor of a magnetically and elastically (by stress frozen-in to the solid component) distorted neutron star. These differences may be enhanced if magnetar-FRB have larger magnetic fields than the known Galactic magnetar-SGR/AXP, implying larger triaxiality of magnetar-FRB moments of inertia. Triaxial neutron stars have been considered in other contexts (Wada & Shimoda, 2024; Shimoda & Wada, 2024).
@@ 2 How Triaxiality May Explain the Three Mysteries?
If a neutron star’s moment of inertia is triaxial by this definition then its rotation is not periodic, but more complex. A triaxial moment of inertia tensor results from frozen-in magnetic stress $B^{2}/8\pi=4\times 10^{28}B_{15}^{2}$ dyne/cm2 and frozen-in elastic stress ${\cal O}(10^{29}$ dyne/cm2) (Horowitz & Kadau, 2009). These may be of similar magnitude and each makes the moment of inertia tensor triaxial. Rotational distortion relaxes to axisymmetry on a hydrodynamic or viscous time scale, is likely smaller than the other distortions for rotational periods $\gtrsim 1\,$s.
@@
    Luo et al. (2025); Zhang & Hu (2025) suggested that in repeating FRB the magnetic and spin axes are aligned, while Kumar, Lu & Bhattacharya (2017); Beniamini & Kumar (2025) suggested that FRB radiation is emitted along open field lines. This leads to a possible explanation (Katz, 2025a) of the difference between repeating and apparently non-repeating FRB: Apparently non-repeating FRB have triaxial moment tensors, so that their magnetic poles and directions of radiation wander pseudo-randomly across much of the sky. FRB that are are only minimally tri-axial have their magnetic and rotational axes aligned (dissipation reduces them to their lowest energy state with oblate spheriodal intertia tensors with magnetic, emission and rotation axes aligned); if fortuitously these point to the observer we see them as repeating FRB, while otherwise they are not observable at all.
@@
The first two of these phenomena might be attributed to “plasma weather” (intrinsic stochasticity of emission) in a magnetar’s magnetosphere, but with no explanatory or predictive theory other hypotheses should be considered. “Plasma weather” cannot explain the absence of the periodicity expected for a rotating neutron star with a spheroidal (axisymmetric) moment of inertia tensor because the observable “weather” would be modulated at the rotation period (as is Solar activity).
@@ 3 Euler Equations
Solutions of these equations when all moments of inertia are unequal are complicated and aperiodic (Landau & Lifschitz, 1976). To make order of magnitude estimates, replace the time derivatives by the algebraic factor $1/\tau$, where $\tau$ is a characteristic time scale, the fractions by $\Delta I/I$, a measure of the asymmetry of the moment of inertia tensor, and assume $I_{1}\approx I_{2}\approx I_{3}$ to high accuracy, as the case for a dynamically nearly spherical body, but that all the $|I_{i}-I_{j}|$ ($i\neq j$) are of the same order of magnitude.
We define a single parameter that measures the degree of triaxiality:
${\cal T}={(I_{1}-I_{2})^{2}(I_{1}-I_{3})^{2}(I_{2}-I_{3})^{2}\over[(I_{1}-I_{0})^{2}+(I_{2}-I_{0})^{2}+(I_{3}-I_{0})^{2}]^{3}},$ (2)
where the mean moment
$I_{0}={I_{1}+I_{2}+I_{3}\over 3}.$ (3)
Then
$0\leq{\cal T}\leq{1\over 2}.$ (4)
If the moment tensor is spheroidal (oblate or prolate) two of the $I_{i}$ are equal and ${\cal T}=0$. Note that $\cal T$} may approach its maximum value even if the moment tensor is nearly isotropic (all the $I_{i}$ are nearly equal), as expected for a slowly rotating magnetar. For example, consider
[…]
Then ${\cal T}\to 1/2$ as either $\epsilon\to 0$ or $\epsilon\to I_{0}$.
3.1 Orders of Magnitude
Solutions of these equations when all moments of inertia are unequal are complicated and aperiodic (Landau & Lifschitz, 1976). To make order of magnitude estimates, replace the time derivatives by the algebraic factor $1/\tau$, where $\tau$ is a characteristic time scale, the fractions by $\Delta I/I$, a measure of the asymmetry of the moment of inertia tensor, and assume $I_{1}\approx I_{2}\approx I_{3}$ to high accuracy, as the case for a dynamically nearly spherical body, but that all the $|I_{i}-I_{j}|$ ($i\neq j$) are of the same order of magnitude.
@@ 3.2 Consequences
If the misalignment is comparable to the beam width then the probability of detecting a burst depends on how close to the line of sight is the beam axis because the radiated intensity decreases with increasing deviation. This varies aperiodically if the deviations from an isotropic moment tensor are triaxial (comparable in magnitude, however small). If the observer remains close to the beam axis the probability of detecting a burst depends on how far he is from the axis. This may explain the failure to find periodicity in the rate of detection of bursts in active repeating FRB Li et al. (2021); Xu et al. (2021); Katz (2022, 2025b).
3.3 Detection of Gravitational Waves From Rotating Neutron Stars
Eq. 7 may also be applied to spinning neutron stars that have been suggested as sources of detectable gravitational waves. Candidates include known fast radio pulsars (Crab and Vela) as well as neutron star X-ray sources like Sco X-1 hypothesized (Galaudage et al., 2022) to be spun-up by accretion.
In order to emit gravitational waves they must not be axisymmetric. That need not, in principle, require them to be triaxial; for example, a dynamically prolate spheroid spinning about a major axis has a radiating gravitational quadrupole moment. However, that geometry is contrived and implausible. Such a neutron star is in a high energy state and dissipation would relax it to a nearly oblate configuration. If an oblate spheroid, it would not radiate gravitational waves.
Triaxial deviations (resulting from magnetic or elastic stress) from an oblate spheroid of a spinning neutron star would be small, likely ${\cal O}(10^{-6}I_{0})$, with ${\cal T}={\cal O}(10^{-6}GM/(8\Omega_{rot}^{2}R^{3}))\sim 10^{-3}$ for the Crab pulsar, where $\Omega_{rot}$ is its angular rotation rate. Without knowing the initial conditions of $\omega_{1},\omega_{2},\omega_{3}$} it is impossible to estimate $\omega_{2}$, $\omega_{3}$ and ${\dot{\omega}}_{1}$, where $\omega_{1}$ refers to the present rotational axis (assumed to be close to the axis of greatest moment of inertia), but the possibility that triaxiality could thwart the coherent integration required to detect a weak periodic gravitational strain should be considered. This may be tested for pulsars by coherent integration of the radio signal; if the radio signal integrates coherently then so will the gravitational signal. Some asteroids tumble, implying not only that they are triaxial but that dissipation has been so slow that their angular momenta have not relaxed to their axes of greatest moment of inertia.